OpenStudy (anonymous):

simplify (1/8)-(7/12)(15)

OpenStudy (lalaly):

\[\frac{7}{12} \times 15= \frac{35}{4}=\frac{70}{8}\]\[\frac{1}{8}-\frac{70}{8}=\frac{-69}{8}\]